# Maximizing the Maximum Degree in Ordered Nearest Neighbor Graphs

### v2

@@ 5 Concluding remarks
Recall from the second half of Theorem 1.1 that there exists a set of $n$ points on the line such that for every order, the indegree of each point in the corresponding Nearest Neighbor Graph is at most $\lceil\log{n}\rceil$. Unlike the linear case, we do not know if this construction is close to being tight for larger dimensions or abstract metric spaces.
Note that both in the definition of Nearest Neighbor Graph corresponding to a metric space and in our proof of Theorem 1.3, we didn’t really use the triangle inequality, and thus we could argue in a more general setting of semimetric spaces. Moreover, the exact distances between the points are also not important, since only their relative order determines the Nearest Neighbor Graph. Furthermore, we know from [2] that any order of the distances on an $n$-element set can be realized in the $(n-1)$-dimensional Euclidean space. Hence, a sufficiently strong improvement upon the result of Theorem 1.2 could strengthen the lower bound from Theorem 1.3 as well.
Observe that our proof of Theorem 1.3 works without any changes if instead of a red clique, we could guarantee only a red ‘backward star’, which is a much sparser hypergraph. (A backward star is a 3-uniform hypergraph similar to a forward star, where the only difference is that the common vertex of all the hyperedges has not the smallest, but the largest index.) We are unaware of any variant of Theorem 4.1 that guarantees a backward star in one color or a forward star in (one of) the other color(s) for a much smaller $n$. Note that the proof of Theorem 4.1 works for any coloring of the complete hypergraph, while colorings that come from genuine metric spaces satisfy additional constraints due to transitivity [^1] For instance, in the notation of Section 4, it is not possible that two triples $\{1,2,3\}$ and $\{1,3,4\}$ are blue, while $\{1,2,4\}$ is green, because otherwise we would have $v_{1}v_{2}<v_{1}v_{3}<v_{1}v_{4}<v_{1}v_{2}$, a contradiction..
@@
Problem 1
For an $n$-element metric space $P$ and $p\in P$, let $d(p)$ be the maximum indegree of $p$ in the ordered Nearest Neighbor Graph over all $n!$ possible orders. Can $\sum_{p}2^{-d(p)}$ be larger than $1$?
On the one hand, note that if the sum $\sum_{p}2^{-d(p)}$ is at most one, then $d(p)\geq\lceil\log{n}\rceil$ for at least one $p\in P$, which would strongly improve upon the result of Theorem 1.3. On the other hand, if $\sum_{p}2^{-d(p)}>1$ for some $P$, then all the points of an appropriate ‘iterated blow-up’ $Q$ of $P$ satisfy $d(q)<c\log{|Q|}$ for some $c<1$ and all $q\in Q$, and this would improve the aforementioned upper bound from the second half of Theorem 1.1.
We managed to verify the inequality $\sum_{p}2^{-d(p)}\leq 1$ for all $n\leq 5$ by computer search. Moreover, it is not hard to check that this sum equals exactly $1$ for the metric spaces constructed in the second half of the proof of Theorem 1.1, and that these are not the only examples. However, computer experiments suggest that such metric spaces become rare as $n$ grows. This can be tentatively considered as evidence towards a negative resolution of Problem 1.
{credits}
5.0.1 Acknowledgements
The authors would like to thank the organizers of the Focused Week on Geometric Spanners (Oct 23 – Oct 29, 2023) at the Erdős Center, Budapest, where this joint work began. Research partially supported by ERC Advanced Grant ‘GeoScape’ No. 882971 and by the Erdős Center. Research was also partially supported by the Overseas Research Fellowship of IIIT-Delhi, India.

### v3

@@ 5 Concluding remarks
Recall from the second half of Theorem 1 that there exists a set of $n$ points on the line such that for every order, the indegree of each point in the corresponding Nearest Neighbor Graph is at most $\lceil\log{n}\rceil$. Unlike the linear case, we do not know if this construction is close to being tight for larger dimensions or abstract metric spaces.
Note that both in the definition of Nearest Neighbor Graph corresponding to a metric space and in our proof of Theorem 3, we didn’t really use the triangle inequality, and thus we could argue in a more general setting of semimetric spaces. Moreover, the exact distances between the points are also not important, since only their relative order determines the Nearest Neighbor Graph. Furthermore, we know from [2] that any order of the distances on an $n$-element set can be realized in the (n1)(n-1)-dimensional Euclidean space. Hence, a sufficiently strong improvement upon the result of Theorem 2 could strengthen the lower bound from Theorem 3 as well.
Observe that our proof of Theorem 3 works without any changes if instead of a red clique, we could guarantee only a red ‘backward star’, which is a much sparser hypergraph. (A backward star is a 3-uniform hypergraph similar to a forward star, where the only difference is that the common vertex of all the hyperedges has not the smallest, but the largest index.) We are unaware of any variant of Theorem 5 that guarantees a backward star in one color or a forward star in (one of) the other color(s) for a much smaller $n$. Note that the proof of Theorem 5 works for any coloring of the complete hypergraph, while colorings that come from metric spaces satisfy additional constraints due to transitivity [^1] For instance, in the notation of Section 4, it is not possible that two triples $\{1,2,3\}$ and $\{1,3,4\}$ are blue, while $\{1,2,4\}$ is green, because otherwise we would have $v_{1}v_{2}<v_{1}v_{3}<v_{1}v_{4}<v_{1}v_{2}$, a contradiction..
@@
Problem 1.
For an $n$-element metric space $V$ and $v\in V$, let $d(v)$ be the maximum indegree of $v$ in the ordered Nearest Neighbor Graph over all $n!$ possible orders. Can $\sum_{v}2^{-d(v)}$ be larger than $1$?
On the one hand, note that if the sum $\sum_{v}2^{-d(v)}$ is at most one, then $d(v)\geq\lceil\log{n}\rceil$ for at least one $v\in V$, which would strongly improve upon the result of Theorem 3. On the other hand, if $\sum_{v}2^{-d(v)}>1$ for some $V$, then all the points of an appropriate ‘iterated blow-up’ $Q$ of $V$ satisfy $d(q)<c\log{|Q|}$ for some $c<1$ and all $q\in Q$, and this would improve the aforementioned upper bound from the second half of Theorem 1.
We managed to verify the inequality $\sum_{v}2^{-d(v)}\leq 1$ for all $n\leq 5$ by computer search. Moreover, it is not hard to check that this sum equals exactly $1$ for the metric spaces constructed in the second half of the proof of Theorem 1, and that these are not the only examples. However, computer experiments suggest that such metric spaces become rare as $n$ grows. This can be tentatively considered as evidence towards a negative resolution of Problem 1.
Acknowledgments.
The authors thank the organizers of the Focused Week on Geometric Spanners (Oct 23 – Oct 29, 2023) at the Erdős Center, Budapest, where this joint work began. Research partially supported by ERC Advanced Grant ‘GeoScape’ No. 882971 and by the Erdős Center and by the Ministry of Innovation and Technology NRDI Office within the framework of the Artificial Intelligence National Laboratory (RRF-2.3.1-21-2022-00004). Research was also partially supported by the Overseas Research Fellowship of IIIT-Delhi, India.